Circular No. 7149

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

POSSIBLE SUPERNOVA IN NGC 3198 W. D. Li, University of California at Berkeley, on behalf of the Lick

Observatory Supernova Search (cf. IAUC 6627, 7126), reports the discovery on unfiltered CCD images taken on Apr. 20.2 UT with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of a variable star (mag ∼ 17.8) near NGC 3198. Inspection of earlier images taken by KAIT reveals that the object was clearly present on Apr. 15.2 (mag ∼ 18.4) and 19.2 (mag ∼ 18.0). The new object is located at α = 10^{h}19^{m}46^{s}.81, δ = +45^{o}31′35′′.0 (equinox 2000.0), which is ∼ 85′′.8 west and 83′′.8 south of the nucleus of NGC 3198. KAIT images of the same field on Apr. 3.2 (limiting mag ∼ 17.5), Mar. 29.2 (limiting mag ∼ 17.5), and Mar. 24.2 (limiting mag ∼ 19.0) showed nothing at the position of the new object. SN 1966J, one of the prototype-Ib supernovae, also occurred in NGC 3198 and had a peak magnitude of B = 11.2.

SUPERNOVA 1999bu IN NGC 3786 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1999bu, obtained on Apr. 18.2 UT by P. Berlind with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), reveals it to be a type-Ic supernova, a few weeks after maximum light. The spectrum features a red continuum with many absorption lines, including a prominent Na I D feature, and is quite similar to a spectrum of the type-Ic SN 1987M taken 29 days after maximum light (Filippenko et al. 1990, A.J. 100, 1575). Emission lines from the host galaxy yield a recession velocity of 2750 km/s.

NO SUPERNOVA IN UGC 11093 In addition to the report on IAUC 7146, Y. L. Qiu, Beijing Astronom-

ical Observatory (BAO), reports that a low-dispersion spectrum taken on Apr. 15.79 UT with BAO 2.16-m telescope at Xinglong Station by D. W. Xu, Qiu, C. Leung, and S. Ko shows that the new object reported as SN 1999bs on IAUC 7141 is not a supernova, but a cataclysmic variable star in outburst. The spectrum shows a blue continuum with absorptions of Hβ, Hγ, Hδ, and Na I D lines. The minima of these absorptions are at 486.0, 434.4, 411.3, and 589.0 nm, and their equivalent widths are 0.73, 0.48, 0.77, and 0.27 nm, respectively. Hα is not conspicuous. The spectrum is similar to that of the dwarf nova CY Lyr at maximum.

1999 April 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green